OpenStudy (anonymous):

Find S12 for 1 + 6 + 11 + 16 +…

OpenStudy (mimi_x3):

looks like an arithmetic series..

OpenStudy (mimi_x3):

what have you tried?

OpenStudy (anonymous):

i dont even know how to do it i dont understand it

OpenStudy (mimi_x3):

\[S_n = \frac{n}{2}[2a+(n-1)d]\] this is the formula you use

OpenStudy (anonymous):

oh wow lol can you break that down please

OpenStudy (mimi_x3):

well have you by chance looked at your textbook or something?

OpenStudy (mimi_x3):

the first term is \(a = 1\) and the \(d\) common difference is \(5\)

OpenStudy (anonymous):

i dont have a text book i do it online

OpenStudy (mimi_x3):

then you're finding \(S_{12}\)

OpenStudy (anonymous):

so what am i doing exactly am i adding or subtracting

OpenStudy (mimi_x3):

\[s_{12} = \frac{12}{2} \left[2(1)+(12-1)*5\right] = ...\] since n = 12, a = 1 , d = 5; i just subbed it in

OpenStudy (anonymous):

would it be 342

OpenStudy (mimi_x3):

well if you got \(342\) then i assume that you're right